A uniform magnetic field B exists in a direction perpendicular to the plane of a square frame made of copper wire. The wire has a diameter of 2mm and a total length of 40cm. The magnetic field changes with time at steady rate dB/dt=0.02Ts^-1. What will be the current induced in the frame? (resistivityof copper=1.7x10^-8 ohm m)

Grade:12

1 Answers

Saakshi Kalashetti
5 years ago
I = e/R 
= 1/ R (d¢/dt)
= 1/R. dBA/dt
= A/R.dB/dt
= A/(pl/a).dB/dt
=A .dB/dt.a/pl
= A(dB/dt).πr^2/pl
Substitute the values and you get the answer.
=9.3×10^-2A
=93mA